OpenStudy (anonymous):

showw mee hooww too doo. pleasee(x 9x+7-2x=14

OpenStudy (anonymous):

\[9x + 7 - 2x = 14\] so now: \[9x - 2x = 14 - 7\] Did you understand what i did here?

OpenStudy (anonymous):

7x=7 = 1, yes you were correct.

OpenStudy (anonymous):

why would you change the position of the 7

OpenStudy (anonymous):

to have like terms on the same side of the equation, to make it easier.

OpenStudy (anonymous):

idk i just dont get it.

OpenStudy (anonymous):

i feel you, that can be super frustrating. I'll try to explain the best i can. hold on please:)

OpenStudy (anonymous):

i get it i get it x=1

OpenStudy (anonymous):

9x+ 7 - 2x = 14 okay so first you are trying to solve for x, so in order to make it easier to do that, you'll want to make sure that the numbers with the x are on the same side of the equation, vise versa ( you'll want to make sure that the plain numbers are on the opposite side ) . Which means that

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

like terms thing I was saying. Then you'll solve both sides of the equation so you only have 1 number on each side

OpenStudy (anonymous):

then in order to get x= , you divide the number that is plain, with the number by x, then you'll get the answer!! :) you're welcome.